A vertical steel rod is 3 m long and has a radius of 0.1m. It is
supporting a load of 1 × 105 N. What is the stress in the rod?
1
Expert's answer
2020-05-24T18:01:23-0400

The stress in the rod


"\sigma=\frac{F}{A}=\frac{F}{\pi R^2}\\
=\frac{1\times 10^5\:\rm N}{3.14\times (0.1\:\rm m)^2}=3.2\times 10^6\:\rm Pa"